# Anti-Degeneracy Scheme for Lidar SLAM based on Particle Filter in Geometry Feature-Less Environments

### v1

@@ IV-B2 Enhanced Resampling for Particles
Model Linear Mapping Gaussian Augmentation GPU Acceleration Image Matrix CrossEntropyLoss BCELoss Adam RMSprop Accuracy F1 Score Time(ms)
@@ V Experimental studies and results
We use a laptop computer with an Intel i7-7700HQ CPU and an NVIDIA GeForce GTX 1060 GPU to remotely maneuver an wheeled robot for data collection and validation of anti-degeneracy system. The robot is equipped with an LSlidar M10P lidar, which is powered by an ARM Cortex-A72 64-bit CPU and a Broadcom VideaCore VI GPU, as shown in Fig. 5. In simulations, we adopt a computer with ubuntu 20.04 system with a cpu configuration of 13th Gen Intel Core i7-13700K, GPU of NVIDIA GeForce RTX 2060 super.
Figure 5: Experimental platforms.
@@ V-A1 Dataset Generation
About the dataset, we record the coordinates of the particles in different scenarios, including straight corridor scenes, feature-sparse outdoor scenes, and feature-rich indoor scenes. We compare the ratio $R$ of the variance to the mean of the particle weights with the degeneracy threshold to generate binary labels. The calculation of $R$ is shown in (12).
@@
where $w$ is the particle weight, $\overline{w}$ is the mean of the weights, and $N$ is the total number of particles.
The degeneracy threshold is determined to be 1 by averaging the values of $R$ from a large number of degeneracy samples. Schematics of the scene used to collect the data and several examples of the particle dataset are shown in Fig. 6, where there are 3937 degraded images and 3827 non-degraded images, and the ratio of training set, test set and validation set is 6:2:2.
@@
ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$ ATE $Error$
@@ V-A2 Ablation Experiments
We conduct ablation experiments to evaluate the contributions of different components in our degeneracy detection model. We test various loss functions and optimizers on two model architectures: MobileNetV3 and DD-Model, and measured accuracy, F1 score, and inference time on the test set as shown in Table. I. The results show that MobileNetV3 achieve an accuracy of around 90% and an F1 score of around 0.85, indicating good performance with room for improvement. In contrast, DD-Model achieve an accuracy of 99.3% and an F1 score of 0.972, demonstrating superior degeneracy detection capability and the ability to identify the vast majority of degraded states. Based on these results, we select CrossEntropyLoss and Adam as our hyperparameters.
We also test the impact of GPU acceleration and image matrix on computation time. The results show that these optimizations reduce the computation time from 130ms to 112ms and 45ms respectively, with the fastest speed reaching 10ms, ensuring the real-time performance of degeneracy detection and SLAM. Finally, we verify the contributions of linear mapping and Gaussian augmentation. Experiments show that removing linear mapping and directly feeding the particle coordinates to the model, or removing Gaussian Augmentation, both lead to a decrease in accuracy. This demonstrates that both modules significantly contribute to improving model accuracy.
@@
ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$ ATE $Error$
@@ V-A3 Accuracy of Degeneracy Detection
In the degeneracy detection experiment, we use [12] as the baseline and replace the eigenvalues of the Hessian matrix with those of the covariance matrix $\Sigma$ of particle weights, the calculation is shown in (13)-(14).
$w=\left[w_{1},w_{2},\ldots,w_{N}\right]$ (13)
@@
where $w$ is a vector of particle weights, $\Sigma$ is the covariance matrix of the weights, and $\overline{w}$ is the mean of the weights.
This is because small eigenvalues can reflect the uncertainty in particle estimation, similar to the degeneracy direction indicated by the eigenvalues of the Hessian matrix. Degeneracy states are detected by comparing with the degeneracy threshold defined in [12]. The comparative results on validation set are shown in Table. III. Our method shows significant improvements in all metrics compared to [12], indicating that our detection method has higher accuracy and robustness.
@@ V-B Performance of Anti-Degeneracy Systems in Simulation Environments
We verify the performance of anti-degeneracy SLAM in various environments through simulation experiments. In terms of experimental scenarios, we build a 10m*10m indoor environment, a 140m*1m long straight corridor, a 100m*100m circular corridor, and an open scenario with sparse landmarks, as shown in the first row of Fig. 8. In terms of evaluation indicators, we plot the maps given by Gmapping which is particle filter-based SLAM with 60 particle as well as by the anti-degeneracy SLAM, in rows 2 and 3 of Fig. 8, and also plot the localization trajectories of Gmapping when the number of particles is 60, 45, and 30, respectively, the anti-degeneracy SLAM with 30 particles and the ground truth given by gazebo, as shown in Fig. 7. In addition, the absolute value of the difference between each trajectory and the true value in the x and y dimensions is plotted in Fig. 7, and its mean value ($Error_{x}$, $Error_{y}$) is shown in Table. II. Finally, the accuracy is quantitatively evaluated by absolute trajectory error (ATE) using evo [22].
As can be seen from the maps in Fig. 8, the anti-degeneracy SLAM is able to accurately construct the environmental structure of each scene, while the Gmapping is less effective in constructing maps in degraded environments, with skewed straight corridors in the second scene, corridors that are not closed in the third, and landmarks in the fourth with incorrect relative positions. Meanwhile, the anti-degeneracy SLAM has the highest fit between the localization trajectory and the true value, and $Error_{x}$ and $Error_{y}$ are maintained within 1m indicating that its localization error is small in all dimensions. We find that the ATE metrics of the anti-degeneracy SLAM decreased in all the scenarios, especially in the corridor scenario, the results are the best, with an accuracy improvement of 90.21%. Theoretically, the higher the number of particles, the higher the SLAM accuracy, but the required computational resources will also rise sharply. And the above results show that increasing the number of particles does not significantly improve the robustness of SLAM in degraded scenes, on the contrary, our anti-degeneracy SLAM is able to maintain high localization accuracy with good robustness in degraded as well as non-degraded environments.
@@ V-C Performance of Anti-Degeneracy Systems in Real-World Environments
We also test anti-degeneracy SLAM in real-world environments. For the experimental scenarios, the setup is a 50m*2.5m long corridor, a 60m*3m long corridor, a 34m*24m circular corridor and an outdoor open scenario, as shown in Fig. 10. In terms of evaluation metrics, it is the same as the simulation experiments, in which the trajectory truth values are manually measured.
As shown in lines 2 and 3 of Fig. 10, the anti-degeneracy SLAM constructs accurate maps of the environment. In contrast, the Gmapping produces less accurate maps that exhibit several issues, like long straight corridors appear skewed, circular corridors fail to close properly, and there are errors in the relative position of landmarks in outdoor scenes.
In terms of localization accuracy, it can be seen from Fig. 9 that even if the localization of other SLAM systems is gradually shifted, our proposed SLAM system is able to ensure the accuracy of localization by anti-degeneracy strategy with minimum gap between the true value. And Table. IV demonstrates that the Gmapping has a large ATE in all degraded scenarios and is unable to maintain good localization accuracy. The ATE of the anti-degeneracy SLAM decreases in all scenarios, especially in the outdoor open environment, where the ATE decreases from 5.9426 to 0.8229, which is roughly 86.2%, and the $Error_{x}$ and $Error_{y}$ are both at a lower value level as well. The above experiments show that in degraded scenarios, the sparseness of environmental features leads to insufficient constraints of SLAM, which makes the localization suffer from error accumulation. And anti-degeneracy SLAM can ensure the accuracy of localization and improve the robustness in various environments.
@@ VI Conclusion
This article presents a deep learning-based anti-degeneracy SLAM system. We first propose a linear mapping and Gaussian-based data augmentation to ensure high-quality training datasets. Next, we develop a degeneracy detection model using ResNet and transformer, which effectively classifies particle distributions to determine degeneracy levels. Finally, we design a hierarchical anti-degeneracy strategy that enhances resampling through selection, fusion, and perturbation, providing accurate initial values for pose optimization. The optimization process can adaptively adjust the optimization frequency and sensor trustworthiness based on degeneracy levels, improving global pose estimation and reducing localization drift. We demonstrate the optimality of the model and the contribution of each module in terms of accuracy and computation time through ablation experiments. In addition, simulation experiments and real experiments are conducted to verify the performance of the anti-degeneracy SLAM in various scenarios. The experimental results show that the anti-degeneracy SLAM is able to construct accurate maps for most of the scenarios, and the reduction of ATE indicates the improvement of its localization accuracy with good robustness.

### v2

@@ IV-B2 Enhanced Resampling for Particles
Model
Linear
Mapping
Gaussian
Augmentation
GPU
Acceleration
Image
Matrix
CrossEntropyLoss BCELoss Adam RMSprop Accuracy F1 Score Time(ms)
@@ V Experimental studies and results
We use a laptop with Intel i7-7700HQ CPU and NVIDIA GeForce GTX1060 GPU to remotely maneuver an wheeled robot for data collection and validation. The robot is equipped with LS-M10P lidar, ARM Cortex-A72 64-bit CPU and Broadcom VideaCore VI GPU, as shown in Fig. 5. In simulations, we adopt a computer with 13th Gen Intel Core i7-13700K CPU and NVIDIA GeForce RTX2060 super GPU.
@@ V-A1 Dataset Generation
We record the coordinates of particle in different scenes, as annotations shown in Fig. 5, and label them with binary values according to $R$ in (15). Since particle weight $w$ reflects credibility, a larger variance and smaller mean of $w$ suggest higher degeneracy. We find $R$ is usually greater than 1 in degenerate scenes, so we classify data with $R>1$ as degenerate and data with $R<1$ as non-degenerate. Finally, there are 3937 degenerate images and 3827 non-degenerate images, the ratio of training, test and validation set is 6:2:2.
@@
where $\overline{w}$ is the mean of $w$ and $N$ is particle number.
@@
ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$
@@ V-A2 Ablation Experiments
We evaluate the contribution of different components in DD-Model by ablation experiments. We test various loss functions and optimizers on two architectures: MobileNetV3 and DD-Model. Accuracy, F1 score and inference time on validation set are shown in Table. I. Specifically, we use MobileNetV3_Large and adapt it to degeneracy detection task. The results show that MobileNetV3 achieve about 90% accuracy and 0.85 F1 score, showing good performance but with room for improvement. In contrast, DD-Model achieve 99.3% accuracy and 0.972 F1 score, demonstrating superior detection capability and ability to identify vast majority of degenerate states. Based on these results, we select CrossEntropyLoss and Adam as hyperparameters.
We also test the impact of GPU acceleration and image matrix. The results show these methods reduce inference time from 130ms to 112ms and 45ms respectively, with the fastest speed reaching 10ms, ensuring the real-time performance. Finally, experiments show that directly feeding the coordinates to model without linear mapping or removing Gaussian augmentation reduces accuracy by about 14.1% and 8.1%, respectively. This demonstrates the contribution of linear mapping and Gaussian augmentation to model accuracy.
@@
ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$ ATE $Error_{x}$ $Error_{y}$
@@ V-A3 Accuracy of Degeneracy Detection
In degeneracy detection experiment, we use [16] as baseline and replace the eigenvalues of Hessian matrix with those of the covariance matrix $\Sigma$ of particle weights in (16). This is because these eigenvalues can reflect the uncertainty in particle estimation.
@@
where $w_{i}$ is the weight of $i_{th}$ particle, $\Sigma$ is the covariance matrix of weights and $\overline{w}$ is the mean of weights.
The comparative results on validation set are shown in Table. III. Our method is better to [16] in all metrics, and the accuracy is increased by 19%, which proves the superior detection capability of DD-Model.
@@ V-B Performance of Anti-Degeneracy Systems in Simulation Environments
We set up four simulation environments S1-S4 to carry out the degeneracy optimization experiments. As shown in Fig. 7, S1 is 10m*10m indoor environment, S2 is 140m*1m long straight corridor, S3 is 100m*100m circular corridor and S4 is open scenario with sparse landmarks. We conduct comparative experiments with GMapping [1], cartographer [2] and DN-CSM [14], while our method uses 30 particles. We plot the maps given by GMapping with 60 particles and other methods in Fig. 7, and also plot the localization trajectories of all methods and the ground truth given by gazebo in Fig. 6, including GMapping with 30, 45 and 60 particles. In addition, the absolute value of the difference between each trajectory and true value in x and y dimensions is plotted in Fig. 6, and its mean value ($Error_{x}$, $Error_{y}$) is shown in Table. II. Finally, the accuracy is quantitatively evaluated by absolute trajectory error (ATE) using evo [24]. Specifically, since position accuracy is the primary concern in robot navigation tasks, we only consider the translational component of trajectory in ATE.
It can be seen that all methods have similar performance in S1 and can accurately construct indoor maps. Cartographer, which employs loop closure detection, has the smallest ATE of 0.106. In S2-S4, both cartographer and DN-CSM fail, unable to construct the basic environmental structure, with the map showing varying degrees of stagnation and steering errors in the direction of movement. In S2, the map constructed by GMapping(60) is skewed. Compared to GMapping(30), our method reduces ATE by up to 79.2%, with $Error_{x}$ and $Error_{y}$ maintained around 0.1, enabling the construction of accurate map. In S3, loop closure in the map by GMapping(60) is not closed, whereas our method constructs accurate loop structure, achieving ATE reduction of up to 90.2% and the best fit with GT trajectory. In S4, our method has low $Error_{x}$ and $Error_{y}$, with localization trajectory closest to GT, and ATE reduction of 79.6% compared to GMapping(30). Additionally, we record the average inference time of DD-Model and hierarchical anti-degeneracy strategies, which are 15.6ms and 25.1ms respectively. The above experiments demonstrate that our method has high localization accuracy in various scenarios and can mitigate the negative impact of degenerate environments on SLAM while ensuring real-time performance.
@@ V-C Performance of Anti-Degeneracy Systems in Real-World Environments
We also conduct degeneracy optimization experiments in four real-world environments R1-R4 shown in Fig. 9, R1 is 50m*2.5m long corridor, R2 is 60m*3m long corridor, R3 is 34m*24m circular corridor and R4 is outdoor scenario with sparse landmarks. In terms of evaluation metrics, it is the same as simulations, and GT of trajectory is manually measured.
As shown in Fig. 9, in R1 and R2, cartographer and DN-CSM show movement stagnation and lose localization in the movement direction. Similarly, in R4, they fail to build correct positions between landmarks because of localization drift.
In R1, GMapping(60) has no localization stagnation, but it shows misalignment due to sparse features, resulting in a skewed map. On the contrary, our method build accurate straight corridor structure. As shown in Table. IV, compared with GMapping(30), ATE of our method decreases by about 73.9%, $Error_{x}$ and $Error_{y}$ are lower than 0.03. In R2, GMapping(60) also shows localization deviation. As can be seen from Fig. 8, the error between localization trajectory and GT of our method is the smallest, the error levels of $Error_{x}$ and $Error_{y}$ are the lowest, and ATE is reduced by 40.5% at the highest. In R3, map loops constructed by GMapping(60), cartographer, and DN-CSM all show dislocations to varying degrees. As can be seen from Table. IV, ATE of cartographer is the highest, 9.215, followed by DN-CSM, 5.7387. Compared with these two methods, the ATE of our method is reduced by 94.9% and 91.8% respectively, and finally more accurate localization results and loop maps are obtained. In R4, the relative position of landmarks constructed by GMapping(60) is wrong. Our method not only shows the correct location of landmarks, but also reduces the ATE by about 86.2% compared with GMapping(30) to obtain more accurate localization results. In addition, we record the mean inference time of DD-Model and hierarchical anti-degeneracy strategies, which are 19.3ms and 27.2ms, respectively.
The above experiments show that compared with GMapping, cartographer, and DN-CSM, our method can improve the localization accuracy of SLAM in various environments and ensure real-time performance in real world. Meanwhile, its adaptive features overcome the heavy dependence of cartographer on preset parameters and the dependence of GMapping on the number of particles.
@@ VI Conclusion
This article has presented a deep learning-based anti-degeneracy SLAM. We first proposed linear mapping and Gaussian-based data augmentation to ensure high-quality dataset. Next, we designed DD-Model using ResNet and transformer, which effectively classifies particle distributions to determine degeneracy levels. Finally, we designed a hierarchical anti-degeneracy strategy that enhances resampling, providing accurate initial values for pose optimization. The optimization process can adaptively adjust the frequency and sensor trustworthiness based on degeneracy levels. We demonstrated the optimality of DD-Model and the contribution of each module through ablation experiments. In addition, the results of simulation and real-world experiments showed that the anti-degeneracy SLAM is able to improve global pose estimation and reduce localization drift in various environments. In the future, we will improve the generalization ability of model under conditions of extreme particle numbers.